I need to find the product using the display method for 57 x 24
1
Expert's answer
2012-10-30T11:55:33-0400
First, calculate the average (81 / 2 = 40,5) Then squareit ( = 1640,25) Now subtract largest number from the average (57 - 40,5 = 16,5)
Square your answer ( = 272,25) Subtract the square from the first square
(1640,25- 272,25 = 1368).
And there's your answer: 1368